The measure of an angle is
(8x+8) and the measure of its supplement is (x+10). Find the measure of each angle.

Answer by Cromlix(4381)

Hi there,
Angles are supplementary if they add up to 180 degrees.
Therefore, (8x + 8) + (x + 10) = 180 degrees
=> 8x + 8 + x + 10 = 180
Collect like terms:
=> 8x + x = 180 - 8 - 10
=> 9x + 162
=> x = 18
(8x + 8) = 152 degrees
(x + 10) = 28 degrees.
Hope this helps :-)